# Adversarially Perturbed Precision Matrix Estimation

### v1

@@ Appendix A Proofs
$2\varepsilon Q(\Delta_{\ast}^{1}-\Delta_{\ast}^{2})=\lambda_{\ast}p(|\Delta_{\ast}^{2}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast}^{2})-|\Delta_{\ast}^{1}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast}^{1})),$ (19)
@@
Proposition A.3.
For vector $b\in\mathbb{R}^{d}$ with $b]_{i}\not=0$ for i=1,,di=1,...,d, positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we have that
$\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta=\|b\|_{1}+\varepsilon\operatorname{sign}(b)^{\top}Q\operatorname{sign}(b),$ (20)
@@
When $\varepsilon=0$, the optimization problem $\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta$ has the unique solution at $\operatorname{sign}(b)$.
When $\varepsilon\not=0$, the objective function is convex and the constraint set is a convex compact set. In this way, the solution to the optimization problem is taken at the boundary of the constraint set, i.e., the solution $\Delta^{\ast}$ satisfies that $\Delta^{\ast}]_{i}=\{-1,1\}$. Since $b]_{i}\not=0$, when $\varepsilon$ is very small, flipping the sign of $\Delta^{\ast}$ will decrease the value of the objective function. Thus, the optimal solution is $\operatorname{sign}(b)$ when $\varepsilon\to 0$. ∎
@@
In the probability space $(\bm{P},C,\mathcal{F})$ for variable $\bm{x}$, for each $\bm{\omega}\in C$, it follows from Proposition A.2 and Proposition A.3 that
@@
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
@@
$\Phi_{n}(C)=-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right)|[C]_{kj}|+\delta^{2}\|C\|_{1,1}.$
@@
$+2\eta\sqrt{n}\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right)\left(|[\Sigma^{-1}]_{kj}+\frac{1}{\sqrt{n}}[\bm{U}]_{kj}|-|[\Sigma^{-1}]_{kj}|\right)$
@@
$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|[\bm{x}]_{k}|]\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
@@
$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|[\bm{x}]_{k}|]\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
Since we have $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma)$, then we have $\bm{x}]_{k}\sim\mathcal{N}(\bm{0},[\Sigma]_{kk})$. In this way,
$\mathbb{E}[|[\bm{x}]_{k}|]=\sqrt{\frac{2[\Sigma]_{kk}}{\pi}}.$

### v2

@@ 7 Discussion
$2\varepsilon Q(\Delta_{\ast}^{1}-\Delta_{\ast}^{2})=\lambda_{\ast}p(|\Delta_{\ast}^{2}|^{p-1}\odot\mathrm{sign}(\Delta_{\ast}^{2})-|\Delta_{\ast}^{1}|^{p-1}\odot\mathrm{sign}(\Delta_{\ast}^{1})),$ (41)
@@
Proposition F.2.
For vector $b\in\mathbb{R}^{d}$ with $b]_{i}\not=0$ for $i=1,\ldots,d$, positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we have that
$\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta=\|b\|_{1}+\varepsilon\mathrm{sign}(b)^{\top}Q\mathrm{sign}(b),$ (42)
@@
When $\varepsilon=0$, the optimization problem $\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta$ has the unique solution at $\mathrm{sign}(b)$.
When $\varepsilon\not=0$, the objective function is convex and the constraint set is a convex compact set. In this way, the solution to the optimization problem is taken at the boundary of the constraint set, i.e., the solution $\Delta^{\ast}$ satisfies that $\Delta^{\ast}]_{i}=\{-1,1\}$. Since $b]_{i}\not=0$, when $\varepsilon$ is very small, flipping the sign of $\Delta^{\ast}$ will decrease the value of the objective function. Thus, the optimal solution is $\mathrm{sign}(b)$ when $\varepsilon\to 0$. ∎
@@
In the probability space $(\bm{P},C,\mathcal{F})$ for variable $\bm{x}$, for each $\bm{\omega}\in C$, it follows from Proposition F.1 and Proposition F.2 that
@@
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}(\bm{x}_{i}\bm{x}_{j},\bm{x}_{i^{\prime}}\bm{x}_{j^{\prime}}).$
@@
$\Phi_{n}(C)=-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|\bm{x}^{i}_{k}|\right)|[C]_{kj}|+\delta^{2}\|C\|_{1,1}.$
@@
$+2\eta\sqrt{n}\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|\bm{x}^{i}_{k}|\right)\left(|[\Sigma^{-1}]_{kj}+\frac{1}{\sqrt{n}}\bm{U}_{kj}|-|[\Sigma^{-1}]_{kj}|\right)$
@@
$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|\bm{x}_{k}|]\left(\bm{U}_{jk}\mathrm{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|\bm{U}_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
@@
$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|\bm{x}_{k}|]\left(\bm{U}_{jk}\mathrm{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|\bm{U}_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
Since we have $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma)$, then we have $\bm{x}_{k}\sim\mathcal{N}(\bm{0},\Sigma_{kk})$. In this way,
$\mathbb{E}[|\bm{x}_{k}|]=\sqrt{\frac{2\Sigma_{kk}}{\pi}}.$